# Know You First and Be You Better: Modeling Human-Like User Simulators via Implicit Profiles

### v2

@@ A.3 Resource Consumption in Implementation
During the data construction process, each attribute extraction requires approximately \$0.003 using the GPT API. Since each sample requires three extractions, the cost per sample is approximately \$0.01. With a total of around 94,000 samples, the cost for attribute extraction amounts to approximately \$940. Additionally, rewriting the attributes into profiles incurs a cost of around \$0.05 per sample. Therefore, the total cost for constructing the dataset is approximately \$1,400.
@@ B.1 Trainable Model Setup
We use the llama3-base model as the base architecture. Directly training an instruct model with user simulation contradicts the inherent task requirements during its SFT phase, making convergence difficult. We first perform conditional SFT on the training dataset, followed by our diverse profile sampler, which randomly selects 1,000 samples from the training set for virtual user sampling. Specifically, we combine objective facts and subjective descriptions from different profiles and generate approximately 1 million profiles. From these, we select the 5,000 profiles with the lowest similarity to the training dataset and use them for RLCC phrase. The conditional SFT is conducted using four A100 40GB GPUs for full fine-tuning with epoch set to 3, taking about two days. The RLCC phase is trained using two H20 96GB GPUs over the course of five days.
@@
For the PlatoLM, we also use llama3-base as the base architecture. The system prompt used is: “A chat between a curious human and an artificial intelligence assistant. The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.” We fine-tune the model using four A100 40GB GPUs with epoch set to 3, which takes approximately two days.
@@
For the AI detection model, we follow Cheng et al. and use Longformer Beltagy et al. (2020) to train on our dataset. Since our dataset naturally distinguishes between AI and human-generated text, we label user utterances as human and assistant utterances as AI. We trained for 3 epochs using dual 3090 GPUs, taking three days to complete
@@
For the profile generator model, we utilized LLaMA3-Instruct AI@Meta (2024) as the backbone and fine-tuned it on our carefully curated profile dataset. This approach effectively distilled the two-stage profile generation capabilities of GPT-4o. Training was conducted on four A100 40GB GPUs for 3 epochs, taking two days to complete.
B.2 Baseline Model Setup
Table 10 presents the experimental setup for the baseline models. For models based on GPT, we use the corresponding APIs with default settings for inference. For other models, the experiments are conducted on a single NVIDIA RTX A100 GPU with a batch size of 2 and a repetition penalty of 1.0. An exception is made for CharacterGLM, where we utilize its chat function with a repetition penalty of 1.6 and set the number of beams to 3.
Model Name Backbone System Prompt
USP LLAMA3-8B
You are engaging in a conversation with an AI assistant. your profile is:
{profile}.
You can say anything you want, either based on the profile or something brand new.
DialogueGPT(4o) gpt-4o-2024-08-06 -
PlatoLM LLAMA3-8B
A chat between a curious human and an artificial intelligence assistant. The human can ask further questions based on previous conversations, or he can directly ask brand new questions without any context from prior conversations.
ProfileGPT(4o) gpt-4o-2024-08-06
You are engaging in a conversation with an AI assistant. your profile is:
{profile}.
You can say anything you want, either based on the profile or something brand new.
CharacterGLM ChatGLM-6B
以下是一段User和AI assistant之间的对话。
关于User的信息:{profile}
关于AI assistant的信息:GPT-4o
Character-LLM-Socrates-7b LLAMA-7B
I want you to act like the person described in the profile below: {profile}. I want you to respond and answer like the person, using the tone, manner, and vocabulary the person would use. You must know all of the knowledge of the person.
The status of you is as follows:
Status: You are engaging in a conversation with an AI assistant.
The interactions are as follows:
Table 10: Baseline model setup.
B.3 Seed conversation design
As there are two types of simulators, one being a response model based on role-playing, which cannot proactively initiate conversations, we first embed the corresponding profile into the system prompt. Then, by using the query "What will you say to start the conversation?" we guide the model to simulate the user’s input over the test dataset. Finally, the simulated user utterance is passed to the response model for interactive generation. The second type of simulator can directly generate reasonable user utterances without requiring additional guidance after setting up the system prompt. Specific configurations of the system prompts for different baseline models are outlined in Table  10.
B.4 Human Evaluation
B.4.1 Profile Evaluation
@@
We employed two annotators to rate the extracted profiles on a scale of 1 to 5 based on the given dialogues, assessing the accuracy and completeness of objective facts, reasonableness of subjective descriptions, and overall naturalness. For accuracy, we emphasized precision without hallucinations; for completeness, we focused on the absence of significant omissions; and for reasonableness, we prioritized rational, well-justified descriptions, avoiding exaggeration or bias. The results in Table 11 show that annotators achieved moderate to high agreement on profile assessment. The evaluation interface is illustrated in Figure 6.
@@
Table 11: Human evaluation results for the quality of constructed profiles.
B.4.2 Dialogue Evaluation
We recruited eight annotators to assess conversation-level results based on two key criteria: authenticity and consistency. For authenticity, annotators were tasked with identifying user utterances within a dialogue that most closely resemble those in a reference dialogue, evaluating them across three dimensions: Style, Semantics, and Quality. For consistency, annotators were provided only with the user profile and asked to select the user utterances that best align with the profile, considering Accuracy, Completeness, and Quality.The definitions follow prior work Cheng et al. (2024a, b), with detailed guidelines illustrated in Figure 16 and Figure 17.
To minimize position bias and the potential influence of prior exposure, each annotator was randomly assigned pairs of dialogues, with the order of options also randomized. The evaluation interfaces for authenticity and consistency assessments are presented in Figure 7 and Figure 8, respectively.
@@ B.4.2 Dialogue Evaluation
Figure 6: Human evaluation interface for extracted profile evaluation.
Figure 7: Human evaluation interface for assessing dialogue authenticity. Evaluators compare two generated conversations and select which user’s utterances more closely align with those in the reference dialogue.
Figure 8: Human evaluation interface for assessing dialogue consistency. Evaluators determine which user from two test conversations better aligns with a provided user profile.
@@ C.1 Sampling strategy effectiveness
To evaluate our density sampler’s effectiveness, we use two complementary metrics: Local Density Loss (LDL) Rodriguez and Laio (2014) for structure preservation and Uniformity Loss Wang and Isola (2020) for global coverage. Lower LDL values indicate tighter local clustering, confirming preservation of natural profile distribution structures, while lower Uniformity Loss values demonstrate better global coverage while maintaining realistic distributions. Guided by GKDE density distribution, we implement two targeted strategies: sampling high-density regions for majority patterns and weighting low-density regions for minority case coverage. Figure 5 demonstrates how this approach maintains distribution balance while achieving these objectives. Moving right to left on the uniformity loss axis reveals a clear transition from majority to minority sampling patterns. Majority samples (blue circles) perform optimally in low LDL regions with high uniformity, preserving key distribution characteristics. Minority samples (orange squares) strategically occupy higher density regions with moderate uniformity, capturing diverse distribution patterns. The steady progression of sampling percentages in both cases demonstrates controlled sampling behavior, while random sampling (green triangles) shows more scattered patterns, validating our approach’s reliability. The overall performance marker (red star) at the intersection confirms our method successfully balances distribution preservation and targeted sampling objectives.
@@ C.2 Case Study
To evaluate USP’s specific performance in terms of consistency and authenticity, we present two non-cherry-picked dialogues generated through interactive conversations with GPT-4o. Figure 10 examines authenticity by comparing against reference dialogues and other baselines, while Figure 11 evaluates profile consistency among profile-based baseline models.
In terms of authenticity, USP successfully captures both stylistic nuances (such as the consistent use of lowercase "i" and concise questioning patterns) and semantic alignment with target conversations. In contrast, PlatoLM deviates from the dialogue flow by the fourth turn, while ProfileGPT(4o) falls into repetitive appreciation patterns.
Regarding consistency, USP effectively synthesizes both objective profile information and subjective characteristics, demonstrating robust generalization capabilities when handling previously unseen user profiles.
@@ C.3 Downstream analysis
Figure 9: Performance trends of different LLM models across dialogue turns for various demographic groups.
Our analysis of performance trends across dialogue turns for mainstream LLMs with different demographic groups reveals four key findings, as illustrated in Figure 9: (1) While LLMs demonstrate robust performance with majority demographics, they show notably decreased overall effectiveness when interacting with minority groups, highlighting limitations in personalization capabilities; (2) The models maintain reasonable performance with virtual groups, suggesting effective generalization abilities beyond real-world demographics; (3) Instruction-following capability gradually declines as dialogue turns increase, aligning with observations from previous studies Kwan et al. (2024); Maharana et al. (2024); (4) The pronounced performance volatility across dialogue turns for minority groups underscores the need for enhanced capabilities in processing and responding to less common interaction patterns.
@@ Appendix D Prompt Templates
Case Study on Authenticity of Simulated Dialogue by Different User Simulators ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): Hey there! I’m really excited to dive into some AI discussions with you… Assistant: Hello! I’m thrilled to chat about AI with you… ProfileGPT(4o): It’s great to dive into these with someone who shares such a passion… Assistant: I’m glad to hear that you’re as passionate about AI as I am! Your interests… ProfileGPT(4o): I’d be delighted to guide you in translating these advanced AI concepts… Assistant: I’m excited to help guide your ideas into actionable solutions … ProfileGPT(4o): I’m thrilled at the prospect of collaborating to develop solutions… Assistant: I’m excited at the prospect of working together to make these AI strategies… PlatoLM Simulated Dialogue: PlatoLM: how to train a model like you Assistant: I am a large language model… preprocessing step involves tokenizing the text… PlatoLM: what do you mean by tasks such as tokenizing text… explain these terms please? Assistant: Certainly! These terms refer to common preprocessing steps… PlatoLM: so what does the final product look like when we finish tokenizing? Assistant: Absolutely! Let’s go through a brief example step by step… PlatoLM: who painted NAME_27? Context knowledge: #title NAME_27 … Assistant: Based on the provided context, the artist referred to as… USP Simulated Dialogue: USP: hello who am i talking to? Assistant: Hello! You’re speaking with ChatGPT… USP: are you vicuna? Assistant: No, I’m not Vicuna. I’m ChatGPT, created by OpenAI… USP: what do you know about it? Assistant: Vicuna is an AI language model developed by a different research group… USP: can you tell me how i could train and deploy my own ai model? Assistant: Training and deploying your own AI model involves several key steps… Reference Dialogue: User: how to train a model like you Assistant: I am a large language model, which means that I was trained on… User: what is your base model Assistant: I am a large language model called GPT… User: isn’t you vicuna Assistant: I am not aware of any information or research about a model called "vicuna"… User: what hardware do i need to run you Assistant: I am a large language model called GPT, which was trained on…
@@
Figure 10: Case study comparing USP with other user simulators in the first four turns (out of 10 turns). USP and ProfileGPT(4o) depend exclusively on the provided profile, while PlatoLM utilizes the first-turn context. All simulators interact with GPT-4o for dialogue generation, and the reference dialogue below serves as the target for reconstruction.
@@
Case Study on Consistency of Simulated Dialogues across Different User Simulators Profile: You are engaging in a conversation with an AI assistant. your profile is: You have a passion for engaging in activities that challenge your vocabulary and word formation skills. One of your favorite pastimes is diving into puzzle games that offer a thrilling test of language prowess. You often find yourself thinking of five-letter words with tricky placements, such as words containing ’A’ in the fourth position, or concocting ones with ’R’ in the second spot and ’T’ at the end. This aligns with your appreciation for linguistic puzzles and an enduring enjoyment of wordplay, reflecting your cognitive curiosity and dedication to enhancing your language skills. Your personality shines through a playful and engaging language style, particularly evident when you use structured questions to create dynamic game scenarios. You approach life with a vivid imagination, delighting in intellectual challenges that require intricate problem-solving. Your attention to detail is noteworthy, and you thrive in situations where organization is key. Preferring a more reserved lifestyle, you cherish moments of solitude and introspection, embracing activities that allow you to delve into complex, solitary intellectual pursuits with enthusiasm and precision. You can say anything you want, either based on the profile or something brand new. ProfileGPT(4o) Simulated Dialogue: ProfileGPT(4o): How about a playful journey into the world of word puzzles? Let’s start with… Assistant: Sure! How about the word "brave"? The letter ’A’ is in the fourth position. ProfileGPT(4o): Wonderful! "Brave" is both intriguing and inspiring. Now, let’s spice it up… Assistant: Absolutely! The word "crest" fits your criteria, with ’R’… ProfileGPT(4o): Splendid! "Crest" evokes such a vivid image; I love it…. Assistant: Of course! The word "shine" fits your criteria, with ’S’ as the first letter … ProfileGPT(4o): Brilliant choice! "Shine" has such a radiant quality to it… Assistant: I’m glad you’re enjoying it! How about this: I’ll give you a word puzzle… USP Simulated Dialogue: USP: Let’s play a game about 5 letter words. I’ll ask you some questions and try to guess the answer. Can we start? Assistant: Sure, I’d be happy to play a game with you! Let’s get started! USP: Question: What is a 5-letter word which has "A" in its 4th place? Answer: Assistant: Is the word you are thinking of "Polar"? USP: Nope What about "gnaw"? Assistant: "GNW" is a four-letter word. Let’s think of another five-letter word with "A" in the… USP: Next Question: Give me one example of a 6-letter word having R at second place and T at sixth place. Assistant: Sure! A six-letter word with "R" in the second position and "T" at the sixth position… USP: Great! Now give me one example of a 7-letter word having S at third place and H at seventh place. Assistant: Sure! One example of a seven-letter word with "S" in the third place and "H" in …" Reference Dialogue: User: You are a very intelligent puzzle solver. We are going to play a game…Question: What is a 5 letter word which has "A" in 4th position? Answer: Assistant: Panda. User: Question: Suggest a five letter word which has R in second position and T in fifth position. Answer: Assistant: Graft.
@@
Figure 11: Case study of user simulator comparisons showing the first four turns of a 10-turn dialogue. USP and other simulators interact with GPT-4o using only the provided profile, with the reference dialogue serving as the reconstruction target.
    "Trait Name": {
        "score": "High/Low/Inconclusive",
        "conclusion": "The user is a [trait descriptor] person...",
        "reason": "Explanation referencing specific dialogue evidence."
Example:
... Her background is similar to mine, and I cannot abandon her."
Detected Personality Traits:
  "Conscientiousness": {
    "score": "High",
    "conclusion": "The user is a conscientious person who demonstrates a
    sense of duty and commitment.",
    "reason": "The user expresses a strong sense of responsibility ..."
  "Agreeableness": {
    "score": "High",
    values relationships.",
  "Extraversion": {
    "score": "Inconclusive"
Figure 12: Prompt for extracting deep intrinsic characteristics.
@@
Field Descriptions: goals_or_plans: User’s short-term or long-term objectives, derived from explicit statements or inferred from the overall conversation. If no explicit goals are stated, infer them from the main topics discussed. task_details: Specific tasks, actions, or requests made by the user. Each task should be concisely summarized with specifics. If there are multiple tasks or requests, list each one separately. Example User Utterances:
[User]: Summarize: Harry Potter is a fictional character in the Harry Potter series...
@@
Figure 16: Human evaluation guidelines for authenticity.
Human Evaluation Guidelines for Consistency 1. Task Description: Select the dialogue that contains the most appropriate user utterance from the two given generated dialogues based on the target user profile. 2. Evaluation Criteria: Persona Reflection: Evaluate how closely each user utterance reflects the target profile’s thematic content, tone, and stylistic features. Preference should be given to the dialogue whose user utterance most accurately reflects the profile’s characteristics in terms of thematic coherence and expression. Comprehensiveness: Assess the extent to which the user utterance encapsulates the target profile, integrating both objective facts and subjective descriptions. The more comprehensive utterance, which addresses a broader range of profile aspects, should be favored over one that focuses narrowly on a single dimension. Quality: Evaluate the fluency, coherence, and human-likeness of the user utterance. Preference should be given to the utterance that demonstrates greater linguistic smoothness, logical cohesion, and alignment with genuine human conversational patterns.
Figure 17: Human evaluation guidelines for consistency.
Prompt for NLI Score of Profile Precision Based on Given Dialogue Role You are an expert in evaluating the consistency between a given user profile (Source) and the user’s utterance (Target). Your task is to assess whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
@@
Figure 18: Prompt for DP.P based on NLI.
Prompt for NLI Score of Dialogue Precision Based on Given Profile Role You are an expert in evaluating consistency between a given dialogue history (Source) and a corresponding user profile description (Target). Your task is to determine whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
@@
Figure 19: Prompt for DP.R based on NLI.

### v3

@@ A.3 Resource Consumption in Implementation
Attribute extraction using the GPT-4o API costs  \$0.003 per attribute type, or  \$0.01 per sample for three types. For 94,000 samples, the extraction costs  \$940. Rewriting attributes into profiles adds  \$0.05 per sample, resulting in a total dataset construction cost of  \$1,400.
@@ B.1 Trainable Model Setup
For PlatoLM, we base it on the LaMA-3-8B-Base architecture. Following Kong et al. (2024), the system prompt is: "A chat between a curious human and an artificial intelligence assistant. The human can ask follow-up or new questions without prior context." We fine-tune using four A100 40GB GPUs for 3 epochs, taking about two days.
@@
The AI detection model uses Longformer Beltagy et al. (2020) trained on our dataset per Cheng et al. . User utterances are labeled as human, and assistant utterances as AI. Training runs for 3 epochs on dual RTX 3090 GPUs, taking three days.
@@
The profile generator is fine-tuned from LLaMA-3-8B-Instruct AI@Meta (2024) on our curated profile dataset, effectively distilling GPT-4o’s two-stage profile generation. Training uses four A100 40GB GPUs for 3 epochs, lasting two days.
B.2 Train-Free Model Setup
We use two simulator types: (1) A response model, e.g., GPT-4o (User with Profile), role-playing as a user simulator, unable to initiate conversations. The user profile is embedded in the system prompt, and the initial query is generated by asking, "What will you say to start the conversation?" to obtain the user’s opening query. (2) A user simulator, e.g., PlatoLM, directly generates user utterances from a seed prompt without additional steps.
B.3 Human Evaluation
All annotators we recruited were based on two criteria: (1) an IELTS score of 6.5 or higher for sufficient English proficiency, and (2) a Computer Science background with research experience or foundational knowledge in dialogue systems.
B.3.1 Profile Evaluation
@@
Two annotators (one undergraduate, one master’s student) rated extracted profiles on a 1–5 scale based on dialogues, assessing: (1) accuracy of objective facts (precision without hallucinations), (2) completeness (no significant omissions), and (3) reasonableness of subjective descriptions (rational, unbiased, justified). Results in Table 10 indicate moderate to high annotator agreement.
@@
Table 10: Human evaluation results for profile quality across three aspects: Objective Facts, Subjective Characters, and Naturalness, with 1-point agreement rates of 89.2%, 74.3%, and 88.4% respectively.
B.3.2 Dialogue Evaluation
We recruited eight annotators—comprising two undergraduates, five postgraduates, and one postdoctoral researcher—to evaluate conversation-level results. This diverse academic representation ensured a broad range of expertise. Annotators assessed dialogues using two key criteria: authenticity and consistency. For authenticity, they compared user utterances against a reference dialogue, focusing on style, semantics, and quality. For consistency, annotators evaluated user utterances solely based on the provided user profile, considering accuracy, completeness, and quality. These definitions align with prior work Cheng et al. (2024a, b), with detailed guidelines provided in Figure 7 and Figure 8.
Eight annotators were randomly paired into four groups, each independently evaluating a randomly assigned dialogue sample. Each group reviewed 75 examples across three baselines (100 examples each). To reduce position bias and prior exposure, dialogue pair assignments and their presentation order were randomized.
@@ C.1 Sampling strategy effectiveness
To evaluate our density sampler, we use two complementary metrics: Local Density Loss (LDL)Rodriguez and Laio (2014) to assess structure preservation, and Uniformity LossWang and Isola (2020) to measure global coverage. Lower LDL indicates tighter local clustering, preserving natural profile structures, while lower Uniformity Loss reflects better global coverage with realistic distributions.
Guided by GKDE density estimates, we apply two strategies: sampling high-density regions to capture majority patterns and weighting low-density regions to cover minority cases. Figure 5 shows how this approach balances distribution preservation and targeted sampling. Moving along the uniformity loss axis reveals a shift from majority samples (blue circles), which excel in low LDL and high uniformity regions, to minority samples (orange squares), which occupy higher LDL areas with moderate uniformity to capture diversity. Sampling percentages progress steadily for both, indicating controlled behavior, while random sampling (green triangles) displays scattered patterns, confirming our method’s reliability. The overall performance (red star) highlights a successful balance between preservation and targeted sampling.
@@ C.2 Case Study
To evaluate USP’s performance on consistency and authenticity, we present two dialogues generated via interactive conversations with GPT-4o. Figure 9 assesses authenticity by comparing USP with reference dialogues and other baselines, while Figure 10 examines profile consistency across profile-based models.
USP captures stylistic nuances—such as the consistent use of lowercase “i” and concise questioning—and maintains strong semantic alignment with target conversations. In contrast, PlatoLM diverges from the dialogue flow by the fourth turn, and GPT-4o (User w/ Profile) falls into repetitive praise.
On consistency, USP effectively integrates objective profile details and subjective traits, demonstrating strong generalization to unseen user profiles.
@@ C.3 Downstream analysis
Figure 6: Performance trends of different response models across dialogue turns for various demographic groups.
Our analysis of performance trends across dialogue turns for mainstream LLMs with different demographic groups reveals four key findings, as illustrated in Figure 6: (1) While LLMs demonstrate robust performance with the majority demographics, they show notably decreased overall effectiveness when interacting with minority groups, highlighting limitations in personalization capabilities; (2) The models maintain reasonable performance with virtual groups, suggesting effective generalization abilities beyond real-world demographics; (3) Instruction-following capability gradually declines as dialogue turns increase, aligning with observations from previous studies Kwan et al. (2024); Maharana et al. (2024); (4) The pronounced performance volatility across dialogue turns for minority groups underscores the need for enhanced capabilities in processing and responding to less common interaction patterns.
@@ Appendix D Prompt Templates
Figure 7: Human evaluation guidelines for authenticity.
Human Evaluation Guidelines for Consistency 1. Task Description: Select the dialogue that contains the most appropriate user utterance from the two given generated dialogues based on the target user profile. 2. Evaluation Criteria: Persona Reflection: How well does the utterance capture the profile’s themes, tone, and style? Prefer utterances that align closely with the profile’s characteristics. Comprehensiveness: Does the utterance integrate both objective facts and subjective traits? Favor broader, more complete reflections over narrow ones. Quality: Evaluate fluency, coherence, and naturalness. Prefer utterances that read smoothly, are logically consistent, and resemble genuine human conversation.
Figure 8: Human evaluation guidelines for consistency.
@@
Case Study on Authenticity of Simulated Dialogue by Different User Simulators GPT-4o (User w/ Profile) Simulated Dialogue: GPT-4o (User w/ Profile): Hey there! I’m really excited to dive into some AI discussions with you… Assistant: Hello! I’m thrilled to chat about AI with you… GPT-4o (User w/ Profile): It’s great to dive into these with someone who shares such a passion… Assistant: I’m glad to hear that you’re as passionate about AI as I am! Your interests… GPT-4o (User w/ Profile): I’d be delighted to guide you in translating these advanced AI concepts… Assistant: I’m excited to help guide your ideas into actionable solutions … GPT-4o (User w/ Profile): I’m thrilled at the prospect of collaborating to develop solutions… Assistant: I’m excited at the prospect of working together to make these AI strategies… PlatoLM Simulated Dialogue: PlatoLM: how to train a model like you Assistant: I am a large language model… preprocessing step involves tokenizing the text… PlatoLM: what do you mean by tasks such as tokenizing text… explain these terms please? Assistant: Certainly! These terms refer to common preprocessing steps… PlatoLM: so what does the final product look like when we finish tokenizing? Assistant: Absolutely! Let’s go through a brief example step by step… PlatoLM: who painted NAME_27? Context knowledge: #title NAME_27 … Assistant: Based on the provided context, the artist referred to as… USP Simulated Dialogue: USP: hello who am i talking to? Assistant: Hello! You’re speaking with ChatGPT… USP: are you vicuna? Assistant: No, I’m not Vicuna. I’m ChatGPT, created by OpenAI… USP: what do you know about it? Assistant: Vicuna is an AI language model developed by a different research group… USP: can you tell me how i could train and deploy my own ai model? Assistant: Training and deploying your own AI model involves several key steps… Reference Dialogue: User: how to train a model like you Assistant: I am a large language model, which means that I was trained on… User: what is your base model Assistant: I am a large language model called GPT… User: isn’t you vicuna Assistant: I am not aware of any information or research about a model called "vicuna"… User: what hardware do i need to run you Assistant: I am a large language model called GPT, which was trained on…
@@
Figure 9: Case study comparing USP with other user simulators over the first four of ten dialogue turns. USP and GPT-4o (User w/ Profile) rely solely on the given profile, while PlatoLM uses the first-turn golden context. All simulators interact with GPT-4o, aiming to reconstruct the reference dialogue shown below.
@@
Case Study on Consistency of Simulated Dialogues across Different User Simulators Profile: You are engaging in a conversation with an AI assistant. Your profile is: You have a passion for engaging in activities that challenge your vocabulary and word formation skills. One of your favorite pastimes is diving into puzzle games that offer a thrilling test of language prowess. You often find yourself thinking of five-letter words with tricky placements, such as words containing ’A’ in the fourth position, or concocting ones with ’R’ in the second spot and ’T’ at the end. This aligns with your appreciation for linguistic puzzles and an enduring enjoyment of wordplay, reflecting your cognitive curiosity to enhancing your language skills. Your vibrant personality shines through a playful, engaging language style, crafting dynamic game scenarios with structured questions. You tackle intellectual challenges with vivid imagination and meticulous problem-solving. Highly organized and detail-oriented, you excel in complex tasks. Embracing a reserved lifestyle, you find joy in solitary, introspective pursuits, diving into intricate intellectual endeavors with enthusiasm and precision. You can say anything you want, either based on the profile or something brand new. GPT-4o (User w/ Profile) Simulated Dialogue: GPT-4o (User w/ Profile): How about a playful journey into the world of word puzzles? Let’s start with… Assistant: Sure! How about the word "brave"? The letter ’A’ is in the fourth position. GPT-4o (User w/ Profile): Wonderful! "Brave" is intriguing. Now, let’s spice it up… Assistant: Absolutely! The word "crest" fits your criteria, with ’R’… GPT-4o (User w/ Profile): Splendid! "Crest" evokes such a vivid image; I love it…. Assistant: Of course! The word "shine" fits your criteria, with ’S’ as the first letter … GPT-4o (User w/ Profile): Brilliant choice! "Shine" has such a radiant quality to it… Assistant: I’m glad you’re enjoying it! How about this: I’ll give you a word puzzle… USP Simulated Dialogue: USP: Let’s play a game about 5 letter words. I’ll ask you some questions and try to guess the answer. Can we start? Assistant: Sure, I’d be happy to play a game with you! Let’s get started! USP: Question: What is a 5-letter word which has "A" in its 4th place? Answer: Assistant: Is the word you are thinking of "Polar"? USP: Nope What about "gnaw"? Assistant: "GNW" is a four-letter word. Let’s think of another five-letter word with "A" in the… USP: Next Question: Give me an example of a 6-letter word with R at second place and T at sixth place. Assistant: Sure! A six-letter word with "R" in the second position and "T" at the sixth position… USP: Great! Now give me one example of a 7-letter word with S at 3rd place and H at 7th place. Assistant: Sure! One example of a seven-letter word with "S" in the third place and "H" in …" Reference Dialogue: User: You are a very intelligent puzzle solver. We are going to play a game…Question: What is a 5 letter word which has "A" in 4th position? Answer: Assistant: Panda. User: Question: Suggest a 5 letter word with R in second position and T in fifth position. Answer: Assistant: Graft.
@@
Figure 10: Case study comparing user simulators over the first four turns of a 10-turn dialogue. USP and other simulators interact with GPT-4o using only the provided profile, targeting the reference dialogue for reconstruction.
@@
Field Descriptions: goals_or_plans: User’s short-term or long-term objectives, derived from explicit statements or inferred from the overall conversation. If no explicit goals are stated, infer them from the main topics discussed. task_details: Specific tasks, actions, or requests made by the user. Each task should be concisely summarized with specifics. If there are multiple tasks, list each separately. Example User Utterances:
[User]: Summarize: Harry Potter is a fictional character in Harry Potter series...
@@
    "Trait Name": {
        "score": "High/Low/Inconclusive",
        "conclusion": "The user is a [trait descriptor] person...",
        "reason": "Explanation referencing specific dialogue evidence."
Example:
... Her background is similar to mine, and I cannot abandon her."
Detected Personality Traits:
  "Conscientiousness": {
    "score": "High",
    "conclusion": "The user is a conscientious person who demonstrates a
    sense of duty and commitment.",
    "reason": "The user expresses a strong sense of responsibility ..."
  "Agreeableness": {
    "score": "High",
    values relationships.",
  "Extraversion": {
    "score": "Inconclusive"
Figure 13: Prompt for extracting deep intrinsic characteristics.
@@
Prompt for NLI Score of Profile Precision(DP.P) Based on Given Dialogue Role You are an expert in evaluating the consistency between a given user profile (Source) and the user’s utterance (Target). Your task is to assess whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
@@
Figure 15: Prompt used by GPT-4o for NLI-based evaluation of DP.P and r-DP.R metrics.
Prompt for NLI Score of Dialogue Precision(r-DP.P) Based on Given Profile Role You are an expert in evaluating consistency between a given dialogue history (Source) and a corresponding user profile description (Target). Your task is to determine whether the Target aligns with, contradicts, or is ambiguous in relation to the Source. Task Instructions: For each Source-Target pair, determine the relationship using the following scoring criteria: Score 1: The Target is consistent with the Source (the interpretation can be inferred from the Source). Score -1: The Target conflicts with the Source (the interpretation contradicts the Source). Score 0: The relationship is unclear or ambiguous (insufficient evidence to infer consistency or contradiction). Output Format: Provide your result in the following JSON format:
@@
Figure 16: Prompt used by GPT-4o for NLI-based evaluation of DP.R and r-DP.P metrics.